### v2

# Structural Chirality and Natural Optical Activity across the -to‑ Phase Transition in SiO2 and AlPO4 from first-principles

Abstract
Natural optical activity (NOA), the ability of a material to rotate the plane of polarized light, has traditionally been associated with structural chirality. However, this relationship has often been oversimplified, leading to conceptual misunderstandings, particularly when attempts are made to directly correlate structural handedness with optical rotatory power. In reality, the relationship between chirality and NOA is more nuanced: optical activity can arise in both chiral and achiral crystal structures, and the sign of the rotation cannot necessarily be inferred from the handedness of the space group. In this work, we conduct a first-principles investigation of natural optical activity in SiO2 and AlPO4 crystals, focusing on their enantiomorphic structural phase transition from high-symmetry hexagonal ( or ) to low-symmetry trigonal ( or ) space groups. This transition, driven by the condensation of a zone-center phonon mode, reverses the screw axis type given by the space group symbol while leaving the sign of the optical activity unchanged. By following the evolution of the structure and the optical response along the transition pathway, we clarify the microscopic origin of this behavior. We demonstrate that the sense of optical rotation is determined not by the nominal handedness of the screw axis given in the space group symbol, but by the atomic-scale handedness of the most polarizable atoms of the structure.
I Introduction
Natural optical activity (NOA) was first observed in the early 19th century by Arago [1] and Biot [2] when they discovered that certain materials could rotate the plane of polarization of light. Because atomic theory and electromagnetic theory were still nascent, the microscopic cause of this rotation and its broader significance remained unclear at the time. Since then, there have been several attempts to give a microscopic explanation of how NOA emerges in certain materials. Many theories have been developed [3, 4, 5, 6, 7], either applying electromagnetic theory or relying on atomic polarizabilities, as described in Ref. [8]. Currently, NOA is understood as a formal representation of the initial spatial dispersion of the macroscopic permittivity tensor as described by first-order principles [9, 10]. Moreover, it has recently become computationally accessible through first-principles density functional perturbation theory in both molecular and solid states efficiently and conveniently [11].
Finally, from an experimental point of view, NOA manifests itself through optical rotation, which can be measured by the rotation angle of the orientation of the plane of polarization about the optical axis of linearly polarized light as it travels through certain materials [12, 3, 13, 14, 15, 16].
While it is a well-known fact that NOA is frequently linked with structural chirality, it is currently understood that chirality is a sufficient, albeit not necessary, condition for NOA to occur, given that achiral materials can also display NOA characteristics. [17, 18, 13]. This historical association led to the widespread belief that the handedness of a material’s space group directly determines the handedness of its optical rotation [19, 20]. Although this correlation may hold in many cases, it is more nuanced: the direction of optical rotation does not necessarily follow the handedness given in the space group symbol. [21, 22]
Measurements of optical activity are a common practice for organic chiral molecules (e.g., sugars or tartaric acid) and, although comparatively weaker, remain measurable in inorganic crystals such as low-quartz and berlinite [21]. Quartz-crystal (SiO2) adopts a trigonal structure (enantiomorphic space groups, 21 (152) or 21 (154)) at room temperature (“-quartz”) and transforms sharply to hexagonal “-quartz” (enantiomorphic space groups, (181) or (180) respectively) at 573 °C; both polymorphs are chiral despite being built from achiral SiO4 tetrahedra. Berlinite (AlPO4), the structural analogue of quartz, similarly undergoes an transition at approximately 586 °C, passing from trigonal () symmetry to hexagonal () without disrupting the network of corner-sharing tetrahedra [23, 24, 25]. In such transitions, the sense of the principal screw axis of the space group symbol reverses sign; however, the sense of the optical activity of the crystal remains unchanged [26, 22]. Initially, it might appear counterintuitive, but the direction of rotation for the chain composed of the most polarizable atoms (which is unchanged throughout the transition) actually determines the direction of optical activity, as shown by Glazer and Stadnicka. [21]. The arrangement of small- and large-channel helices endows quartz with a chiral bivector [27, 28]. This explains why the optical activity is preserved despite the reversal of the principal screw axis in these crystal structures.
In this work, we aim to carry out an in-depth first-principles investigation of this structural transition and, following the workflow presented in Ref. [29], we show that it can be associated with the progressive condensation of a zone-center phonon mode consistent with symmetry analysis of rotations of SiO4 units [30]. By following the evolution of the crystal structure from the high-symmetry (HS) to the low-symmetry (LS) enantiomorphic phase, we track how both the screw axis and the natural optical activity evolve along the transition path, providing microscopic insight into the interplay between chirality, lattice dynamics, and optical rotation.
II Computational details
For all the calculations, we converged the structural data of the parent structure with the Abinit code [31] (version 9.10.5) with the PBEsol exchange-correlation functional. We used the plane wave-pseudopotential approach with optimized norm-conserving pseudopotentials as taken from the PseudoDojo server (v.4) [32, 33] and an energy cutoff of Hartrees. The -mesh sampling employed for the calculations was 666 for the structural relaxations as well as for the computation of the NOA. A value of Har/Bohr was employed on the forces to stop the structural relaxations and a value of Ha was used for the electronic residual self-consistent cycle stop. Phonon calculations were computed through the density functional perturbation theory framework as implemented in abinit [34, 31].
For the calculation of the optical activity, we focus on the low-frequency limit of , which tends to a constant as , where is the rotatory power and the reduced Planck constant. For both SiO2 and AlPO4, we assume that the incident light propagates along the optic axis, which is considered to be parallel to the Cartesian direction. Under these conditions, , where is the speed of light and is a component of the natural optical activity tensor in the zero frequency limit. While numerous methodologies and strategies for computing the NOA in solids exist [4, 5, 35, 7, 36, 37], this study utilizes the latest linear-response approach grounded in density-functional perturbation theory developed by Zabalo and Stengel [11] and incorporated into abinit.
III Results
III.1 Structural transition of SiO2 and AlPO4
Let us start by analyzing the high-symmetry phase of SiO2, which crystallizes in the enantiomorphic (or ) space group. For simplicity, we restrict our analysis to the structure, as the methodology for the other space group is analogous and yields identical conclusions.
This structure is characterized by a network of corner-linked SiO4 tetrahedra, arranged about a principal screw axis aligned with the crystallographic -axis as shown in Fig. 1. Noticeably, the crystallization process fixes the chirality into the (or ) space group and the associated properties.
Recently, an algorithm has been proposed to identify chiral displacive transitions from achiral to chiral space groups [29], following the works of Fava et al. [38] for K3NiO2. Applying this workflow [29] to SiO2 and AlPO4, we find that a displacive phase transition from an achiral to a chiral space group is not allowed by pseudosymmetry. In other words, no achiral phase lies sufficiently close to the chiral phases of SiO2 or AlPO4 to permit a displacive transition.
However, symmetry considerations predict that the high symmetry right-handed screw-axis structure can be linked to the low-symmetry 21 left-handed screw-axis structure [24, 25] through the irrep of the zone center [39, 30] (or similarly between the and the space groups). This distortion removes symmetry, leaving a screw as shown by the shaded hexagon and double triangles in Fig. 1. This means that those two space groups can form a high-symmetry/low-symmetry pairs for SiO2 or AlPO4.
From our density functional theory (DFT) calculations, we confirm the presence of an unstable phonon mode at the point in the phase of SiO2, as referenced by [40]. By checking the corresponding phonon eigenvector symmetry, we found that it has the irrep, i.e. in agreement with the symmetry analysis which links the right-handed screw of the high-symmetry phase and the left-handed screw of the low-symmetry phase. Such a mode is neither Raman nor infrared active, i.e. it is silent as reported some time ago through hyper-Raman [41] or inelastic neutron scattering [42]. Hence, these results go in the direction of the displacive character of the transition [43], helping clarify previous discussions on its nature [44]. The calculated structural data of the converged structure computed with a -mesh of 666 can be found in Table 1. From our density functional perturbation theory (DFPT) calculations, we found that the unstable imaginary frequency of the mode has an amplitude of cm-1. The condensation and relaxation of the eigenvector gives an energy difference between the high and low temperature phases of meV per 9 atoms within the unit cell.
| Atom | x | y | z |
| Si (3d) | 0.0000 | 0.5000 | 0.3333 |
| O (6i) | 0.2080 | 0.4160 | 0.5 |
| Atom | x | y | z |
| Si (3a) | 0.0000 | 0.4689 | 0.6666 |
| O (6c) | 0.1442 | 0.7308 | 0.4515 |
A symmetry adapted mode analysis using the isodistort software[45, 46] of our relaxed low-symmetry phase reveals that the distortion decomposes into and irreps where the amplitude of the mode is marginal (0.03 Å) with respect to the leading mode (1.00 Å) which drives the transition. The mode induces in-plane displacements of the Si atoms and both out-of-plane and in-plane displacements of the O atoms. The displacements are such that the SiO4 tetrahedra rotate as fairly rigid units (and do not distort significantly), as illustrated in Figure 1(c). As shown in Fig. 1(a-b), the in-plane projection of the displacements induces a coordinated counter-clockwise and clockwise movement between the neighboring helical centers. This movement breaks the -fold screw axis at the corners of the unit cells that transform into centers while preserving the -fold screw axis at the interior of the unit cell (see Table. 2).
Moving now to the case of AlPO4 crystal and when studying its 22 phase, we can observe a similar crystal structure to that of the 22 of SiO2 [47, 48] with the exception that the -axis is essentially doubled due to the ordering of Al + P over the atom sites (as discussed below). The crystal structure is shown in Fig. 2. The calculated structural data of the converged AlPO4 22 structure can be found in Table 3. Similarly to the previous SiO2 case, we also obtain an unstable phonon mode at the point of the right-handed high-symmetry phase with an amplitude of cm-1 and the irrep [48]. This mode, hence, gives the link between the right-handed high-symmetry and the left-handed low-symmetry chiral phases, again similarly to the case of SiO2 but with doubled cell along the z direction. The condensation and relaxation of the eigenvector of this unstable mode gives, in this case, an energy gain of meV per 18-atom unit cell, which is comparable to SiO2 when normalized to the number of atoms. The relaxed structure we obtain is fully consistent with previous works [49, 47]. Similar conclusions to those drawn in the previous example can be reached by analyzing the distortion induced by the mode, which causes in-plane displacements of the Al and P atoms and both in-plane and out-of-plane displacements of the O atoms as in the case of the quartz crystal (see Table. 4). The symmetry adapted mode decomposition gives amplitudes of 0.08 Å for the irrep and 1.57 Å for the irrep.
Hence, we found with our calculations that both SiO2 and AlPO4 have an unstable phonon mode in their high-symmetry 22 (or, respectively, 22) chiral phase that drives the phase transition towards the lower temperature chiral structure 21 (or 21). We will now examine the connection between these phase transitions and their optical activity.
| Atom | x | y | z |
| Al (3c) | 0.5000 | 0.0000 | 0.0000 |
| P (3d) | 0.5000 | 0.0000 | 0.5000 |
| O (12k) | 0.1938 | 0.4201 | 0.7560 |
| Atom | x | y | z |
| Al (3a) | 0.4680 | 0.0000 | 0.3333 |
| P (3b) | 0.4682 | 0.0000 | 0.8333 |
| O (6c) | 0.1265 | 0.7104 | 0.2684 |
| O (6c) | 0.1598 | 0.7445 | 0.7821 |
III.2 Evolution of the natural optical activity
To explore the microscopic origin of the optical activity in these systems, we now analyze the evolution of the NOA along the structural transition path from the high-symmetry to the low-symmetry enantiomorphic phases in both SiO2 and AlPO4.
| 5.4973 | ||
| 5.8443 | ||
| 6.2176 | ||
| 6.4082 | ||
| 6.5985 | ||
| 6.9731 | ||
| 7.1562 | ||
| 0.0 | 7.3364 | |
| 0.1 | 7.1562 | |
| 0.2 | 6.9731 | |
| 0.4 | 6.5985 | |
| 0.5 | 6.4082 | |
| 0.6 | 6.2176 | |
| 0.8 | 5.8443 | |
| 1.0 | 5.4973 |
The results are summarized in Fig. 3 and Table 5. The first interesting remark emerging from our calculations is the opposite sign of the NOA in the high-symmetry high-temperature phases of SiO2 and AlPO4, despite sharing the same enantiomorphic space group. Our first-principles calculations predict a positive rotatory power for the and phases of SiO2 for light traveling along its optical axis, in agreement with Ref. [11] (despite the handedness of the screw axis given in this space group symbol) and a negative rotatory power for light traveling along its optical axis for the and phases of AlPO4. The numerical values we obtained for SiO2 are in good agreement with the values encountered in the literature [50] and the change of rotatory sign between SiO2 and AlPO4 is also consistent with a previous report by Glazer and Stadnicka [21]. This finding underscores a crucial concept: the sign of a crystal’s natural optical activity links to its structural chirality, though it might not align with the helical handedness indicated by the space group symbol [21, 51].
This finding cautions against the common assumption that the sign of rotatory power unambiguously reflects the absolute sense of principal screw-axis [52] even in materials with similar space groups.
We can trace and explain the difference in rotatory power by examining the crystal structures of both materials. As mentioned in the previous section and now shown in Fig. 4, the atomic positions of AlPO4 in its setting are very similar to those of the SiO2 crystal in its setting. In fact, when we perform a computational experiment by substituting the Si atoms into the Al and P sites of the AlPO4 structure, and calculate the optical activity, we obtain a rotatory power with the opposite sign to that of SiO2 in its phase and that closely resembles that found for AlPO4 in its phase. Furthermore, upon relaxing the atomic positions in this modified structure, the system naturally evolves into the phase of SiO2 although initially a symmetry is detected by Abinit. Although the one-to-one alternation of Al and P atoms and the non-equivalence of the O atoms induces a right-handed screw axis of the material as discussed in Ref. [21, 47], Fig. 4(d) shows that a non-symmetric left-handed helix similar to that observed in 22 SiO2 is indeed present in the crystal. Although not perfectly symmetric, this chain formed by the most polarizable atoms in the crystal (oxygen sites in this particular case) dominates the optical activity of these phases explaining its rotatory power [21, 51]. Moreover, the observation that substituting Si atoms at the Al and P sites does not significantly alter the rotatory power suggests that these atomic species contribute little to the optical activity, which mainly arises from the more polarizable oxygen atoms.
The second interesting observation is that the condensation of the -distortion decreases the value of the rotatory power in these types of transitions where the screw-axis type is reverted. Indeed, as shown in Fig. 3 we observe a linear decrease of the rotatory power as we condense the distortion. This can be explained by looking at the distortions shown in Fig. 1 and Fig. 2 that tend to disrupt the helical arrangement of the O atoms. One might expect that further increasing the amplitude of the distortion could eventually revert the sign of the optical activity; however, our calculations show that, for instance, an amplitude 50% larger than the optimal value, recovers rotatory power values comparable to those of the undistorted high-symmetry phase. Moreover, this increase in distortion amplitude is accompanied by a significant increase in the Landau energy as shown in Fig. 3 (b), indicating that such configurations are physically inaccessible.
IV Conclusion
In this work, we performed an ab-initio analysis of the structural transitions and associated changes in the natural optical activity between the high-symmetry () and low-symmetry () phases of SiO2 and AlPO4. Our first-principles calculations show that both materials undergo a displacive phase transition driven by the softening of a phonon mode at the zone center with symmetry. This mode induces a distortion that reverses the sign of the principal screw-axis of the crystal high-symmetry space group.
Despite this reversal, the sense of the natural optical activity of the materials is preserved throughout the transition. AlPO4 and SiO2 therefore provide a useful illustration of the principle that it is the handedness of the most polarizable atoms in a structure that determines the sign of its natural optical activity [21, 51]. Upon cooling through its phase transition, the screw-sense of the imperfectly symmetric helix of oxygen sites is not changed (although this is not obvious from the space group symmetries of the and phases where the helix handedness given by the space group symbol is reversed), preserving therefore the optical activity of the system.
Interestingly, the closely related structures of AlPO4 and SiO2 also highlights the role of the most polarizable oxide ions. The AlPO4 and SiO2 structures differ due to the ordering of Al and P sites in AlPO4 which doubles the axis of the unit cell. As a result, the space groups of the higher symmetry phases are reversed (high temperature AlPO4 of symmetry has atom sites, including those of the polarizable oxygens, close to those of SiO2 of symmetry). This means that the optical activity sign of AlPO4 is the same as that of SiO2.
Furthermore, our analysis of the -transition pathway indicates a linear decrease in the magnitude of the rotatory power as the mode progresses from the high-symmetry to the low-symmetry phase for both the SiO2 and AlPO4 crystals. This trend suggests that the symmetry-lowering introduced by the distortion reduces the handedness of the structure (as shown in Figure 3). Our results highlight the intricate relationship between structural chirality and optical activity in chiral crystals and caution against simplistic interpretations based solely on space group symmetry.
Acknowledgements.
F.G.O. acknowledges financial support from MSCA-PF 101148906 funded by the European Union and the Fonds de la Recherche Scientifique (FNRS) through the FNRS-CR 1.B.227.25F grant. F.G.-O. and E.B. acknowledge the Fonds de la Recherche Scientifique (FNRS) for financial support, the PDR project CHRYSALID No.40003544 and the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the F.R.S.-FNRS under Grant No. 2.5020.11 as well as computational resources made available on Lucia, the Tier-1 supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement No. 1910247. F.G.-O. and E. B. also acknowledge support from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 964931 (TSAR). We also recognize the support of the West Virginia High Education Policy Commission under the call Research Challenge Grand Program 2022, Award RCG 23-007 and NASA EPSCoR Award 80NSSC22M0173References
- Arago [1811] F. Arago, Mémoire sur la polarisation colorée, Œuvres complètes de François Arago (1811).
- Biot [1812] J. B. Biot, Sur de nouveaux rapports qui existent entre la réflexion et la polarisation de la lumière des corps cristallisés, Mem. Cl. Sci. Math. Phys. Inst. 13, 1 (1812).
- Condon [1937] E. U. Condon, Theories of optical rotatory power, Rev. Mod. Phys. 9, 432 (1937).
- Natori [1975] K. Natori, Band theory of the optical activity of crystals, J. Phys. Soc. Jpn. 39, 1013 (1975).
- Zhong et al. [1993] H. Zhong, Z. H. Levine, D. C. Allan, and J. W. Wilkins, Band-theoretic calculations of the optical-activity tensor of -quartz and trigonal Se, Phys. Rev. B 48, 1384 (1993).
- Barron [2009] L. D. Barron, Molecular Light Scattering and Optical Activity (Cambridge University Press, Cambridge, England, 2009).
- Óscar Pozo Ocaña and Souza [2023] Óscar Pozo Ocaña and I. Souza, Multipole theory of optical spatial dispersion in crystals, SciPost Phys. 14, 118 (2023).
- Kizel’ [1980] V. A. Kizel’, Optical activity and dissymmetry in living systems, Phys. Usp. 23, 277 (1980).
- Landau and Lifshits [1984] L. D. Landau and E. M. Lifshits, Electrodynamics of Continuous Media. Volume 8 : Course of Theoretical Physics, Second edition ed. (Pergamon, 1984).
- Agranovich and Ginzburg [1984] V. M. Agranovich and V. Ginzburg, Crystal optics with spatial dispersion, and excitons (Springer-Verlag, New York, 1984).
- Zabalo and Stengel [2023] A. Zabalo and M. Stengel, Natural optical activity from density-functional perturbation theory, Phys. Rev. Lett. 131, 086902 (2023).
- Jerphagnon and Chemla [1976] J. Jerphagnon and D. S. Chemla, Optical activity of crystals, J. Chem. Phys. 65, 1522 (1976).
- Bousquet et al. [2025] E. Bousquet, M. Fava, Z. Romestan, F. Gómez-Ortiz, E. E. McCabe, and A. H. Romero, Structural chirality and related properties in periodic inorganic solids: review and perspectives, Journal of Physics: Condensed Matter 37, 163004 (2025).
- Nomura [1960] K. C. Nomura, Optical activity in tellurium, Phys. Rev. Lett. 5, 500 (1960).
- Ades and Champness [1975] S. Ades and C. H. Champness, Optical activity of tellurium to 20 µm, J. Opt. Soc. Am. 65, 217 (1975).
- Iwasaki et al. [1972] H. Iwasaki, K. Sugii, N. Niizeki, and H. Toyoda, Switching of optical rotatory power in ferroelectric 5PbO 3GeO2 single crystal, Ferroelectrics 3, 157 (1972).
- Halasyamani and Poeppelmeier [1998] P. S. Halasyamani and K. R. Poeppelmeier, Noncentrosymmetric oxides, Chemistry of Materials 10, 2753 (1998).
- Fecher et al. [2022] G. H. Fecher, J. Kübler, and C. Felser, Chirality in the solid state: Chiral crystal structures in chiral and achiral space groups, Materials 15, 10.3390/ma15175812 (2022).
- Wooster [1953] W. A. Wooster, Physical properties and atomic arrangements in crystals, Reports on Progress in Physics 16, 62 (1953).
- Gómez-Ortiz et al. [2024] F. Gómez-Ortiz, M. Fava, E. E. McCabe, A. H. Romero, and E. Bousquet, Structural chirality measurements and computation of handedness in periodic solids, Phys. Rev. B 110, 174112 (2024).
- Glazer and Stadnicka [1986] A. M. Glazer and K. Stadnicka, On the origin of optical activity in crystal structures, Journal of Applied Crystallography 19, 108 (1986).
- Glazer [2018] A. M. Glazer, Confusion over the description of the quartz structure yet again, Journal of Applied Crystallography 51, 915 (2018).
- Ng and Calvo [1976] H. N. Ng and C. Calvo, X-ray study of the - transformation of berlinite (AlPO4), Canadian Journal of Physics 54, 638 (1976).
- Shapiro et al. [1967] S. M. Shapiro, D. C. O’Shea, and H. Z. Cummins, Raman scattering study of the alpha-beta phase transition in quartz, Phys. Rev. Lett. 19, 361 (1967).
- Grimm and Dorner [1975] H. Grimm and B. Dorner, On the mechanism of the - phase transformation of quartz, Journal of Physics and Chemistry of Solids 36, 407 (1975).
- Donnay and Le Page [1978] J. D. H. Donnay and Y. Le Page, The vicissitudes of the low-quartz crystal setting or the pitfalls of enantiomorphism, Acta Crystallographica Section A 34, 584 (1978).
- Hlinka [2014] J. Hlinka, Eight types of symmetrically distinct vectorlike physical quantities, Phys. Rev. Lett. 113, 165502 (2014).
- Erb and Hlinka [2018] K. C. Erb and J. Hlinka, Symmetry guide to chiroaxial transitions, Phase Transitions 91, 953 (2018).
- Gómez-Ortiz et al. [2025] F. Gómez-Ortiz, A. H. Romero, and E. Bousquet, Pathways to crystal chirality an algorithm to identify new displacive chiral phase transitions (2025), arXiv:2503.13076 [cond-mat.mtrl-sci] .
- Campbell et al. [2018] B. Campbell, C. J. Howard, T. B. Averett, T. A. Whittle, S. Schmid, S. Machlus, C. Yost, and H. T. Stokes, An algebraic approach to cooperative rotations in networks of interconnected rigid units, Acta Crystallographica Section A 74, 408 (2018).
- Gonze et al. [2020] X. Gonze, B. Amadon, G. Antonius, F. Arnardi, L. Baguet, J.-M. Beuken, J. Bieder, F. Bottin, J. Bouchet, E. Bousquet, N. Brouwer, F. Bruneval, G. Brunin, T. Cavignac, J.-B. Charraud, W. Chen, M. Côté, S. Cottenier, J. Denier, G. Geneste, P. Ghosez, M. Giantomassi, Y. Gillet, O. Gingras, D. R. Hamann, G. Hautier, X. He, N. Helbig, N. Holzwarth, Y. Jia, F. Jollet, W. Lafargue-Dit-Hauret, K. Lejaeghere, M. A. Marques, A. Martin, C. Martins, H. P. Miranda, F. Naccarato, K. Persson, G. Petretto, V. Planes, Y. Pouillon, S. Prokhorenko, F. Ricci, G.-M. Rignanese, A. H. Romero, M. M. Schmitt, M. Torrent, M. J. van Setten, B. Van Troeye, M. J. Verstraete, G. Zérah, and J. W. Zwanziger, The Abinitproject: Impact, environment and recent developments, Computer Physics Communications 248, 107042 (2020).
- Hamann [2013] D. Hamann, Optimized norm-conserving vanderbilt pseudopotentials, Physical Review B 88, 085117 (2013).
- Van Setten et al. [2018] M. J. Van Setten, M. Giantomassi, E. Bousquet, M. J. Verstraete, D. R. Hamann, X. Gonze, and G.-M. Rignanese, The pseudodojo: Training and grading a 85 element optimized norm-conserving pseudopotential table, Computer Physics Communications 226, 39 (2018).
- Gonze and Lee [1997] X. Gonze and C. Lee, Dynamical matrices, Born effective charges, dielectric permittivity tensors, and interatomic force constants from density-functional perturbation theory, Physical Review B 55, 10355 (1997).
- Malashevich and Souza [2010] A. Malashevich and I. Souza, Band theory of spatial dispersion in magnetoelectrics, Phys. Rev. B 82, 245118 (2010).
- Wang and Yan [2023] X. Wang and Y. Yan, Optical activity of solids from first principles, Phys. Rev. B 107, 045201 (2023).
- Urru et al. [2025] A. Urru, I. Souza, O. P. Ocaña, S. S. Tsirkin, and D. Vanderbilt, Optical spatial dispersion via wannier interpolation, Phys. Rev. B 112, 045201 (2025).
- Fava et al. [2024] M. Fava, E. McCabe, A. H. Romero, and E. Bousquet, A phonon-driven mechanism for an emergent and reversible chirality in crystals (2024), arXiv:2405.12696 [cond-mat.mtrl-sci] .
- Antao [2016] S. M. Antao, Quartz: structural and thermodynamic analyses across the transition with origin of negative thermal expansion (NTE) in quartz and calcite, Acta Crystallographica Section B 72, 249 (2016).
- Raman and Nedungadi [1940] C. V. Raman and T. M. K. Nedungadi, The - transformation of quartz, Nature 145, 147 (1940).
- Tezuka et al. [1991] Y. Tezuka, S. Shin, and M. Ishigame, Observation of the silent soft phonon in -quartz by means of hyper-raman scattering, Phys. Rev. Lett. 66, 2356 (1991).
- Axe and Shirane [1970] J. D. Axe and G. Shirane, Study of the quartz phase transformation by inelastic neutron scattering, Phys. Rev. B 1, 342 (1970).
- Choudhury and Chaplot [2006] N. Choudhury and S. L. Chaplot, Ab initio studies of phonon softening and high-pressure phase transitions of -quartz , Phys. Rev. B 73, 094304 (2006).
- Dolino et al. [1983] G. Dolino, J.-P. Bachheimer, F. Gervais, and A. F. Wright, La transition - du quartz; le point sur quelques problèmes actuels: Transition ordre-désordre ou displacive, comportement thermodynamique, Bulletin de Minéralogie 106, 267 (1983).
- [45] H. T. Stokes, D. M. Hatch, and B. J. Campbell, ISODISTORT, ISOTROPY software suite, iso.byu.edu.
- Campbell et al. [2006] B. J. Campbell, H. T. Stokes, D. E. Tanner, and D. M. Hatch, ISODISPLACE: a web-based tool for exploring structural distortions, J. Appl. Crystallogr. 39, 607 (2006).
- Schwarzenbach [1966] D. Schwarzenbach, Verfeinerung der struktur der tiefquarz-modifikation von alpo4, Zeitschrift für Kristallographie - Crystalline Materials 123, 161 (1966).
- Muraoka and Kihara [1997] Y. Muraoka and K. Kihara, The temperature dependence of the crystal structure of berlinite, a quartz-type form of AlPO4, Physics and Chemistry of Minerals 24, 243 (1997).
- Thong and Schwarzenbach [1979] N. Thong and D. Schwarzenbach, The use of electric field gradient calculations in charge density refinements. II. Charge density refinement of the low-quartz structure of aluminum phosphate, Acta Crystallographica Section A 35, 658 (1979).
- Jönsson et al. [1996] L. Jönsson, Z. H. Levine, and J. W. Wilkins, Large local-field corrections in optical rotatory power of quartz and selenium, Phys. Rev. Lett. 76, 1372 (1996).
- Glazer and Stadnicka [1989] A. M. Glazer and K. Stadnicka, On the use of the term ‘absolute’ in crystallography, Acta Crystallographica Section A 45, 234 (1989).
- Tanaka et al. [2010] Y. Tanaka, T. Kojima, Y. Takata, A. Chainani, S. W. Lovesey, K. S. Knight, T. Takeuchi, M. Oura, Y. Senba, H. Ohashi, and S. Shin, Determination of structural chirality of berlinite and quartz using resonant x-ray diffraction with circularly polarized x-rays, Phys. Rev. B 81, 144104 (2010).